OpenStudy (anonymous):

Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. -185 -77 -23 85

OpenStudy (tkhunny):

Why not substitute and evaluate?

OpenStudy (anonymous):

Well, \[3^{3} - 2y ^{2} + -3^{4} ?\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Sorry, -2(5)^2

OpenStudy (tkhunny):

Where did the z go? Substitute the value for 'y', too. Get them all!

OpenStudy (tkhunny):

x^3 - 2y^2 - 3x^3 + z^4 x = 3 (3)^3 - 2y^2 - 3(3)^3 + z^4 y = 5 (3)^3 - 2(5)^2 - 3(3)^3 + z^4 z = -3 (3)^3 - 2(5)^2 - 3(3)^3 + (-3)^4 Just like that.

OpenStudy (anonymous):

3^3 = 27

OpenStudy (anonymous):

2(5)^2 = 20

OpenStudy (anonymous):

3(3)^3 = 81?

OpenStudy (anonymous):

-3^4 = 81

OpenStudy (tkhunny):

Not quite. Let's try that again. (3)^3 - 2(5)^2 - 3(3)^3 + (-3)^4 3^3 = 27 -- You got that. 27 - 2(5)^2 - 3(3)^3 + (-3)^4 2(5^2) = 2*25 = 50 27 - 50 - 3(3)^3 + (-3)^4 3(3^3) = 81 -- You got this one. 27 - 50 - 81 + (-3)^4 (-3)^4 = +81 -- Careful with the notation. This MAY be different from -3^4. These two expressions are not ALWAYS the same. 27 - 50 - 81 + 81

OpenStudy (anonymous):

-23?

OpenStudy (tkhunny):

Why do you doubt? :-)

OpenStudy (anonymous):

idk. lol

OpenStudy (anonymous):

You were a lot of help, thanks.